OpenStudy (gabbyalicorn):

Which circles have an area of 16π in^2? Choose all answers that are correct. A. a circle with a radius of 2 in. B. a circle with a radius of 4 in. C. a circle with a diameter of 8 in. D. a circle with a diameter of 16 in.

OpenStudy (gabbyalicorn):

i think C B and A

OpenStudy (gabbyalicorn):

@KendrickLamar2014 @misty1212

OpenStudy (gabbyalicorn):

@JackofallTradez @JoannaBlackwelder

OpenStudy (jackofalltradez):

Actually, it would be B and C, because it's (pi)r^2 D(Diameter) = 2r 4^2(pi) = 16(pi)^2

OpenStudy (gabbyalicorn):

Okay, Thank you!

OpenStudy (jackofalltradez):

No problem ^_^